Nitrogen pentoxide reacts with nitric oxide in the
Answer #1


i. let kf and kb be forward and backward rate constants
              kf
N2O5 ==== NO2 + NO3
              kb
ii. NO3 + NO --> 2 NO2           (rate constant k2)


In these steps, NO3 is intermediate. You have


consumption rate of NO3 = k2 [NO3] [NO]

production rate of NO3 = kf [N2O5]

Applying the steady-state assumption gives:

k2 [NO3] [NO] = kf [N2O5]


[NO3] = kf [N2O5]/k2[NO] --- 1

d[NO2]/dt = 2* k2 [NO3][NO] ---- 2


Substituting (1) in (2) and then in (3) gives

d[NO2]/dt = 2*k2(kf [N2O5]/k2[NO])[NO]


d[NO2]/dt = K [N2O5]--- 3

where k = 2*Kf